19. You are evaluating two mutually exclusive projects, A and B, with equal investment of $5,000, required rate of return of 8%, and following cash flows in years 1-3 (for project A) and 1-5 (for project B) Year 1 Year 2 Year 3 Year 4 Year 5 Project A 18001700 3200 Project B 1200 1280 1350 1400 2200 Which project will you prefer and why? A. Project A since it reports higher NPV of 806.5 B. Project B since it reports higher NPV of 806.5 C. Project A since it reports higher EAA of 257.81 D. Project B since it reports higher EAA of 201.99
Answer #1

In case of projects with unequal life, the right meaure to compare the projects is not NPV but equivalent annual benefit

NPV of project A = 1,800*PVF(8%, 1 year) + 1,700*PVF(8%, 2 years) + 3,200*PVF(8%, 3 years)-5,000

= 1,800* 0.926+1700*0.857+3200*0.794-5,000

=$664.5

EAA = NPV/PVAF(8%, 3 years)

= $257.81

NPV of project B = 1200*0.926+1280*0.857+1350*0.794+1400*0.735+2200*0.681 - 5,000

=$806.5

EAA = NPV/PVAF(8%, 5 years)

=$201.99

Since EAA of project A is higher, it should be selected

The correct answer is C.